Parth (parthkohli):

Given: \(4x^2 + y^2 = 40\) \(xy = 6\) Find \(2x + y\).

Parth (parthkohli):

Is this is a start? x = 6/y

ganeshie8 (ganeshie8):

(2x+y)^2 = (4x^2 + y^2)+2xy 2x+y = ?

Parth (parthkohli):

Yeah, that's what exactly is given in the book. How do you derive \((2x + y)^2\)?

OpenStudy (anonymous):

That will make it lengthy parth..

Parth (parthkohli):

I mean how did you get that \((2x + y)^2\)?

ganeshie8 (ganeshie8):

thats the start point... we have a^2, b^2 and ab terms and asking to find a+b

OpenStudy (unklerhaukus):

but there are four solution for (x,y)

OpenStudy (anonymous):

(2x+y)^2-4xy=40 \[2x+y=\sqrt{40+24}\] \[2x+y=\sqrt{64}=8\]

OpenStudy (anonymous):

\[(2x + y)^2 = 4x^2 + y^2 + 4xy \implies 64\] \[(2x + y) = \pm 8\]

OpenStudy (anonymous):

it should be \[\pm 8\]

Parth (parthkohli):

Hmm I am still understanding it.

Parth (parthkohli):

\( \color{Black}{\Rightarrow (2x + y)^2 = 4x^2 +4xy + y^2 }\)

OpenStudy (anonymous):

Parth, first equation gives you : \(4x^2 + y^2 = 40\) and second \(xy = 6\) Just plug these in the formula: \[(2x + y)^2 = 4x^2 + y^2 + 4xy\]

OpenStudy (unklerhaukus):



Parth (parthkohli):

Oh wait. Getting it.

OpenStudy (anonymous):

\[(2x + y)^2 = \color{green}{(4x^2 + y^2)} + 4 \color{green}{(xy)} \implies (40) + 4(6)\]

Parth (parthkohli):

\( \color{Black}{\Rightarrow (2x + y)^2 = 64}\) \( \color{Black}{\Rightarrow 2x + y = \pm 8}\)

Parth (parthkohli):

Yep

OpenStudy (anonymous):

Yes now you got it right..

Parth (parthkohli):

Thank you!

ganeshie8 (ganeshie8):

but we should get 4 values.. as per the pix of @UnkleRhaukus

OpenStudy (anonymous):

Welcome dear..

ganeshie8 (ganeshie8):

scratch that... me confused :\

OpenStudy (unklerhaukus):